OpenStudy (anonymous):

What is the standard form equation of the line passing through (5, –3) with a slope of –3? 3x + y = 12 3x + y = –12 3x – y = 12 3x – y = –12

OpenStudy (anonymous):

i think some one else is helping you on this

hero (hero):

Given: (x,y) = (5,3) m = -3 Find b: 1. Write slope-intercept formula y = mx + b 2. Substitute given values: (-3) = (-3)(5) + b 3. Solve for b: -3 = -15 + b -3 + 15 = b 12 = b 4. Write the equation of the line (include values of m and b only) y = -3x + 12 5. Add 3x to both sides: 3x + y = 12

hero (hero):

(x,y) = (5,-3) my bad

OpenStudy (anonymous):

Thanks man! :)

hero (hero):

You can thank me with a medal